OpenStudy (anonymous):

Using a directrix of y=5 and a focus of (4,1), what quadratic function is created?

OpenStudy (anonymous):

a. f(x)= 1/4 (x-4)^2 -3 b. f(x)= 1/8(x+4)^2 -3 c. f(x)= -1/8 (x-4)^2 +3 d. f(x)= -1/4 (x+4)^2 -3

OpenStudy (anonymous):

did you try to make a sketch?